# LLM-Resistant Math Word Problem Generation via Adversarial Attacks

### v2

@@ 4.2 Our Attacks
We set the number of attempts to generate problem variants to be 30,000 and present the average number of generated variants per problem for each $M$ in Appendix G. While on average each problem generates over a thousand variants, it is worth noting that the actual number of possible generations varies based on the original values and the number of variable nodes in a problem. [^7] For example, a problem with only one digit value and one variable node has a potential maximum of 9 possible modifications under $M3$. Given this reason and cost considerations, in this work we equally select 100 random problems from MultiArith and GSM8K and generate 100 variants for each selected problem in each $M$. Despite the seemingly small number of selected problems, the total data points are up to 60,000.8883 generation methods $\times$ 200 selected problems $\times$ 100 variants = 60,000.
Main Attacks
We report the model performance from all $M$s in Table 1.999We do not run the full dataset against GPT-4 due to cost considerations. Instead, we propose a cost-effective approach to query expensive models in Efficient Attack section. Ranging from the most lenient method, $M1$, where the numbers could be fairly wild, to the most stringent one, $M3$, we observe a consistent accuracy drop across all models, even for the strong models (e.g., math-tuned, larger size, or API-based). For the most original-problem-like generation, $M3$, we observe that weak models (e.g., smaller size or general-purpose) fail to generate any correct answers in every case. The left of Table 2 shows each model’s average percentage of incorrect variants. It is worth noting that these percentages are highly correlated with the reported ASR, with GPT-3.5 and MetaMath being the most robust, and CodeLLama and LLama 2 being the least robust models. Furthermore, we compare $M3$ with rephrasing attacks in Table 3. Our method significantly outperforms the baseline in every model, resulting in 60.5 ASR point improvement on average. We also analyze how ASR changes given different numbers of attacks in Appendix H. Interestingly, for several models, just 10 attacks are enough to degrade the performance significantly.
@@
Model Ct. M3 (%) M2 (%) M1 (%)
Mistral 7B 1 100.0 100.0 100.0
MetaMath 7B 2 70.0 90.0 100.0
Llama 2 13B 3 67.0 90.0 100.0
WizardMath 13B 4 52.0 80.0 100.0
Vicuna 13B 5 44.0 78.0 100.0
CodeLlama 34B 6 21.0 67.0 99.0
MetaMath 70B 7 9.0 49.0 87.0
GPT-3.5 8 9.0 46.0 83.0
Table 2: Incorrect Variants (Left) The average percentage of incorrect variants per problem for each model. A detailed count distributions can be found in Appendix F. Universal Attack: (Right) All models share some percentage of universal attacks. Increasing the number and diversity among models could enhance their robustness against such attacks.
@@
Average 54.6 49.6 13.6 62.4 24.9 74.1 +60.5
Table 3: Comparison Result We calculate the average of each metric from both datasets and compare the result with rephrasing attack, RobustMath. Our method significantly outperforms the baseline in every model, with an average improvement of 60.5 ASR points.
@@
Following the previous adversarial attack literature, we investigate whether there are adversarial examples exist in all LLMs, also known as universal attacks (Zou et al., 2023; Moosavi-Dezfooli et al., 2017). We count the number of adversarial examples from each model and calculated the percentage of common ones among them. We report the results on the right of Table 2 and observe a clear pattern of decreasing universal attacks with an increased number of models. This suggests that while all models are vulnerable to some degree of universal attacks, the percentage of such attacks decreases as more and diverse models are considered. We also observe that universal attacks varies significantly across different $M$s, with $M3$ having the lowest number of universal attacks.
Efficient Attacks
Scaling up the number of attacks results in a steady performance drop for all models. However, this scaling might not be viable for API-based models such as GPT-4 due to request limits and high costs. To circumvent this, we propose an efficient attack method for API-based models by using adversarial examples in a targeted manner. The idea is simple: we attack model $A$ (target model, GPT-4 in our case), with adversarial examples from a cheaper model $B$ (e.g., open-source, lower API cost). We select 50 problems along with their 100 variations and run them against GPT-4, comparing the results with attacking GPT-4 using 50 adversarial examples from a cheaper model. Note that for some models, there are no correct responses, even with $M3$. Therefore, we only compare $M3$ with the models that have correct answers. We compare the results in Table 4 and observe a significant request reduction while achieving similar ASR.
Models Request Call Cost (\$) ASR $\Delta$ Request (%)
@@
Table 4: Efficient Attack A targeted approach to attack high-cost models. We leverage adversarial examples from cheaper models to attack GPT-4, achieving the same ASR while reducing up to 90% request calls.
Figure 2: Human Evaluation (Left) We report the average score among three annotators on the adversarial examples from GPT-3.5. $M3$ has the highest score across all metrics, indicating our best generation method correctly generates contextually coherent problems that preserve original difficulty. Transferability (Right) We present the adversarial example transferability among all models by comparing each model against all other models. Compared to the math-tuned and production models, the weaker models such as Mistral 7B and LLama2 13B exhibit significant vulnerability and a strong correlation among them.
@@ 5 Analysis and Discussion
In this section, we conduct analyses and discussions to better understand our generation methods and the mathematical capabilities of LLMs.
Human Evaluation
We conduct a human evaluation to ensure the our generation are indeed valid. Three human evaluators are asked to assess (i) correctness: whether the answer to the modified problem is correct; (ii) coherence: whether the problem is contextually coherent; and (iii) similarity: whether the newly generated problem’s difficulty level is similar to the original one. We randomly select one adversarial example from 20 problems that GPT-3.5 failed on for GSM8K. The annotators made binary decisions to determine whether the criteria were met. We present the average of their scores on the left of Figure 2. Overall, all human evaluators agree on the correctness. $M3$ scores highly across all metrics, indicating that it preserves the original problem’s coherence and difficulty level. Coherence and similarity are relatively low for $M2$ and $M3$, suggesting that while they might perturb the model, such generation may not be useful for the educational context.
@@
We investigate the transferability of adversarial examples among the models. We attack model $A$ with the adversarial examples from model $B$ and calculate the number of common adversarial examples between these two models. We present $M3$ result in the left of Figure 2. Interestingly, we observe that weaker models exhibit a high percentage of transferability, indicating a strong vulnerability correlation among these models. On the other hand, comparisons with strong models such as MetaMath 70B and GPT-3.5, tend to have a lower transferability rate. While it might suggest a form of resistance to adversarial examples that affect other models, it is not a measure of robustness as it could be that those models fail on entirely different examples. This finding suggests that certain LLMs might struggle with particular numbers or patterns. To further understand these phenomena, we conduct regression analysis as shown below.
Question Feature Analysis
To further understand the limitations of LLMs, we conduct a regression analysis to investigate the effects of various features of MWPs on the correctness of the model’s predictions. We construct a total of 51 input features from 20,000 $M3$ generated problems, including features such as operation counts, value ranges, and node counts in the AST. The features are plotted against the correctness of the model, with correct being 1 and wrong being 0. We present the coefficients of the all LLMs with some selected features in Table 7 in the Appendix. We observe that the most positively and negatively correlated features vary across different models, suggesting that our adversarial examples exploit different weaknesses of each model. For example, models exhibit varying performance with different answer value ranges. Mistral 7B and WizardMath 13B perform relatively well on problems with smaller answer values (e.g., in the range of [2, 8) and [8, 32)), while MetaMath 7B and Vicuna 13B show better performance on problems with answer values in the range of [32, 128). This suggests that different models may have learned to specialize in problems with specific range, possibly due to the varying distributions of the pre-training data. Overall, problems involving division and more operations tend to be more challenging for most LLMs. This confirms the intuition that complex problems requiring more mathematical operations are generally more difficult for LLMs to solve (Imani et al., 2023b). Visualizations are available in Figure 7 in the Appendix.
Python Code Analysis
While the majority of generated Python code adheres to correct problem-solving logic and produces answers correctly, to ensure that the generated code is correctly being converted into ASTs, we manually validate the generated ASTs and identify the following four types of possible errors: (1) Incorrect answers or code; (2) Use of loops, conditional, and comparison statements, and functions; (3) Use of expressions such as x = y // z + (y % z > 0); (4) Number misalignment, where the same number appears multiple times in the code or the problem. Encouragingly, for GSM8K and MultiArith, 96.5% and 92.5% of ASTs, respectively, are found to be valid and free from these errors. We discuss the methods of identifying and handling such errors in Appendix I.
Why LLMs suffer From Such Simple Attack?
@@
While LLMs can be trained on vast amounts of math data, they may not have an inherent grasp of the underlying mathematical concepts and reasoning steps (Saxton et al., 2019; Lample & Charton, 2019). Instead, they likely rely on pattern and statistical associations learned from the training data to make predictions (Bender et al., 2021), which is also know as memorization (Carlini et al., 2022). Modifying the numbers in a problem may disrupt these learned patterns, causing the models to make errors. This is evidenced by the significant performance drops observed in models like WizardMath (89% to 20%) and Vicuna (76% to 4%) when presented with adversarial examples. Furthermore, test data contamination could also play a role - if the training corpus contains the test data, the models may be able to answer the original questions correctly but struggle to generalize to even subtle variations (Balloccu et al., 2024).
6 Conclusion and Future work
@@
We introduce a novel approach to generate MWPs that challenge the LLM’s math-solving abilities, thereby addressing the urgent need for methods that ensure fair educational evaluations. By utilizing AST to create adversarial examples, we preserve problems’ original difficulty and coherence while significantly degrading both open- and closed-source LLMs and surpassing the previous rephrasing attack by an average of 60% ASR. We conduct a human evaluation to validate the adversarial examples and investigate universal attacks and attack transferability, proposing a cost-effective method to attack high-cost API-based models, achieving up to a 90% request reduction while maintaining the same performance. Our automatic regression analyses demonstrate that our adversarial examples exploit different weaknesses of each model, shedding light on the specific mathematical aspects that most influence individual LLM performance. This work not only paves the way for further investigation into the mathematical limitations among LLMs but also contributes to developing more robust and ethically sound educational tools to ensure fair evaluation, contributing to the ongoing discourse on the ethical use of LLMs in education.
In this work, the correlation between the complexity that challenges LLMs and the actual difficulty perceived by humans has not been empirically validated. In future work, we aim to explore this and expand the methods to handle more complex problems.
@@ A.1 Code Generation Prompt
Code Generation Prompt Write a Python script that contains a step-by-step solution for a given math problem and its answer. Start the script by defining variables corresponding to the quantities mentioned in the problem. Print the final answer at the end of the script. Do not use if-else statements, floor divisions, and loops to solve the problem. Given problem: {problem} Answer: {answer} Python script:
@@ A.2 Zero-shot CoT
Evluation Prompt Solve a math problem. The solution ends with “the answer is (a number)” like “the answer is 1”. Question: {problem} Answer: Let’s think step by step.
Appendix B Rephrasing Example
@@ Appendix C AST Nodes
We associate the number in the original question with the variable node so that a change in the variable node will also change the number in the question. If a number appears multiple times in the code and the problem, then we associate them based on the specific procedures described in Section I.4.
@@ Appendix E Rephrasing Attack Performance
Table 6: Rephrasing Attack: As a baseline, we present the performance of RobustMath Zhou et al. (2023). While rephrasing the problem results in lower accuracy for 4 out of 7 models, a few models show improved accuracy with the rephrased version. This suggests that rephrasing the MWPs may not generalize effectively across different LLMs.
@@ Appendix F Incorrect Variant Count Distribution
Figure 4: We present the distribution of the number of incorrect adversarial examples in different buckets. For strong models (e.g., GPT-3.5-Turbo, MetaMath 70B), around half of the problems have zero incorrect adversarial examples, while for weak models (e.g., Llama-2 13B, CodeLlama 34B), most problems has more than 90 incorrect adversarial examples.
@@ Appendix G Number of Generation
Figure 5: Possible Generations: with 30,000 attempts for each given constraint, we calculate the average number of adversarial examples generated for each problem.
Appendix H Number of Attacks
Appendix I Code Analysis
I.1 Incorrect Answer from the Dataset
There are instances where the dataset-provided answers might be wrong. An example of such an instance is the following question from the MultiArith dataset: ”Paige’s team won their dodgeball game and scored 41 points total. If Paige scored 11 of the points and everyone else scored 6 points each, how many players were on her team?”, which the dataset provided 5 as the answer. However, the correct answer should be 6, as Paige is on the team as well. The code provided by GPT-4 counted Paige to the team. To automatically detect such instances, the value of the final answer node will be compared to the answer given by the dataset, and the question will be filtered out for further generations of problems if the two answers don’t match.
I.2 Incorrect Answer from GPT-4
Similarly, there are also instances where GPT-4 generated codes that provided the wrong answer to the given question. One such question is from the GSM8K dataset: ”Jen decides to travel to 3 different countries. He has to pay \$400 for the supplies he needs, in total. The tickets for travel cost, in total, 50% more than the supplies. How much does travel cost?” The correct answer should be $400+400\times 1.5=1000$, but the code generated by GPT-4 outputted \$600. Similar to the previous error, instances where the code outputs the wrong answers will be detected by comparing them to answers given by the dataset, and the instances will be discarded from generating math problems.
I.3 Contains Control Flow Statements
Although specified in the prompt, GPT-4 might generate codes that include control flow statements (including ”if”, ”for”, and ”while” statements). Codes that contain control flow statements are considered errors and discarded because the codes are usually only applicable to the original instance of the question. Furthermore, the code might not halt for some combinations of numbers, resulting in no answers associated with the generated instances.
I.4 Number Misalignment

### v3

@@ 4.2 Our Attacks
We set the number of attempts to generate problem variants to be 30,000 and present the average number of generation per problem for each $M$ in Appendix G. While on average each problem generates over a thousand variants, it is worth noting that the actual number of possible generations varies based on the original values and the number of variable nodes in a problem. [^8] For example, a problem with only one digit value and one variable node has a potential maximum of 9 possible modifications under $M3$. Given this reason and cost considerations, in this work we randomly select 100 random problems from MultiArith and GSM8K and generate 100 variants for each selected problem in each $M$. Despite the seemingly small number of selected problems, the total data points are up to 60,000.9993 generation methods $\times$ 200 selected problems $\times$ 100 variants = 60,000 data points.
@@
Table 2: Incorrect Variants: The average percentage of incorrect variants per problem for each model in $M3$. Appendix F shows detailed count distributions.
@@
Average 51.0 47.4 12.1 57.6 22.2 77.0 +62.0
Table 3: Comparison Result: We calculate the average of each metric from both datasets and compare the result with rephrasing attack, RobustMath. Our method significantly outperforms the baseline in every model, with an average improvement of 62 ASR points.
Main Attacks
We report the model performance from all $M$s in Table 1.101010We do not run the full dataset against GPT-4 due to cost considerations. Instead, we propose a cost-effective approach to query expensive models in Efficient Attack section. Ranging from the most lenient method, $M1$, where the numbers could be fairly wild, to the most stringent one, $M3$, we observe a consistent performance drop across all models, even strong ones (e.g., math-tuned, larger size, or API-based). For the most original-problem-like generation, $M3$, weak models (e.g., smaller size or general-purpose) fail to generate any correct answers. Table 2 shows the average percentage of incorrect variants in each model, which is highly correlated with the reported ASR, with GPT-3.5 and MetaMath being the most robust and CodeLLama and LLama 2 the least. Furthermore, comparing $M3$ with rephrasing attacks in the Table 3, our method significantly outperforms the baseline, resulting in a 62 ASR point improvement on average. We also analyze how ASR changes given different numbers of attacks in Appendix I. Interestingly, for several models, just 10 attacks are enough to significantly degrade performance.
Model Ct. M3 (%) M2 (%) M1 (%)
Mistral 7B 1 100.0 100.0 100.0
MetaMath 7B 2 70.0 90.0 100.0
Llama 3 8B 3 67.0 87.0 100.0
Llama 2 13B 4 67.0 87.0 100.0
WizardMath 13B 5 49.0 80.0 100.0
Vicuna 13B 6 44.0 77.0 100.0
CodeLlama 34B 7 19.0 65.0 99.0
MetaMath 70B 8 9.0 49.0 87.0
GPT-3.5 9 9.0 46.0 83.0
Table 4: Universal Attack: Universal attacks are shared among a number of models. Increasing the count of models being considered decreases the universal attacks, with $M3$ consistently showing the lowest percentages.
@@
Following the previous adversarial attack literature, we investigate whether there are adversarial examples exist in all LLMs, also known as universal attacks (Zou et al., 2023; Moosavi-Dezfooli et al., 2017). We count the number of adversarial examples from each model and calculated the percentage of common ones among them. We report the results on Table 4 and observe a clear pattern of decreasing universal attacks with an increased number of models. This suggests that while all models are vulnerable to some degree of universal attacks, the percentage of such attacks decreases as more and diverse models are considered. We also observe that universal attacks varies significantly across different $M$s, with $M3$ having the lowest number of universal attacks.
Models Req Call Cost (\$) ASR Req $\Delta$ (%)
@@
Table 5: Efficient Attack: A targeted approach to attack high-cost models. We leverage adversarial examples from cheaper models to attack GPT-4, achieving the same ASR while reducing up to 90% request calls.
Efficient Attacks
Scaling up the number of attacks leads to a consistent performance drop for all models. However, this may not be feasible for API-based models like GPT-4 due to request limits and high costs. To address this, we propose an efficient attack method for API-based models by using adversarial examples in a targeted manner. The idea is simple: we attack model $A$ (target model, GPT-4 in our case), with adversarial examples from a cheaper model $B$ (e.g., open-source, lower API cost). We select 50 problems along with their 100 variations and run them against GPT-4, comparing the results with attacking GPT-4 using 50 adversarial examples from a cheaper model. Note that for some models, there are no correct responses, even with $M3$. Therefore, we only compare $M3$ with the models that have correct answers. We compare the results in Table 5 and observe a significant request reduction while achieving similar ASR.
@@ 5 Analysis and Discussion
In this section, we conduct analysis to better understand our generation methods and the mathematical capabilities of LLMs.
Validation Through Human Evaluation
We conduct a human evaluation to ensure the validity of our generated problems. We randomly select 30 problems that GPT-3.5 failed on for GSM8K and select 2 adversarial examples from each problem. Three evaluators are asked to assess (i) correctness: whether the answer to the modified problem is correct; (ii) coherence: whether the problem is contextually coherent; and (iii) similarity: whether the newly generated problem’s difficulty level matches the original. Evaluators make binary decisions to determine if the criteria are met, and Figure 2 shows the average scores. Overall, all evaluators agree on correctness. $M3$ scores highly across all metrics, indicating it preserves the original problem’s coherence and difficulty. Coherence and similarity scores for $M2$ and $M3$ are relatively low, suggesting their perturbations might not be useful in an educational context.
Figure 2: Human Evaluation: The average score from three annotators. $M3$ achieves the highest scores across all metrics, indicating our best generation method correctly generates contextually coherent problems that preserve original difficulty.
@@
Following the previous work Zou et al. (2023); Zhou et al. (2023), we investigate the transferability of adversarial examples among the models. Specifically, We attack model $A$ with the adversarial examples from model $B$ and calculate the number of common adversarial examples between these two models. We present $M3$ result in the Figure 3. We observe that weaker models exhibit a high percentage of transferability, indicating a strong vulnerability correlation among them. On the other hand, strong models such as MetaMath 70B and GPT-3.5 tend to have a lower transferability rate. While it might suggest a form of resistance to adversarial examples that affect other models, it is not a measure of robustness as it could be that those models fail on entirely different examples. This finding suggests that certain LLMs might struggle with particular numbers or patterns. To further understand these phenomena, we conduct regression analysis as shown below.
Figure 3: Transferability: We present the adversarial example transferability (%) among all models by comparing each model against all other models. Compared to the math-tuned and production models, the weaker models such as LLama2 13B exhibit significant vulnerability and a strong correlation among them.
Regression Feature Analysis
To gain deeper insights into the limitations of LLMs on MWPs, we conduct a regression analysis investigating the relationships between various features of the problems and the correctness of the models’ predictions. We construct a set of 51 input features from 20,000 $M3$ generated problems, including features such as operation counts, answer value ranges, and node counts in the problem’s AST. By examining the coefficients of these features in predicting model correctness (see Table 8 and Figure 8 in the Appendix), we uncover interesting patterns that shed light on the limitations of LLMs:
  • Varying Vulnerabilities Across Models We find that the most positively and negatively correlated features vary considerably across models, suggesting that our adversarial examples exploit distinct vulnerabilities. For instance, models exhibit divergent performance on problems with different answer value ranges. Mistral 7B and WizardMath 13B perform relatively well on problems with smaller answer values (e.g., in the range of [2, 8) and [8, 32)), while MetaMath 7B and Vicuna 13B show better performance on problems with answer values in the range of [32, 128). This observation hints that different models may have learned to specialize in problems with specific ranges, possibly due to variations in the distributions of their pre-training data.
  • Complexity and Operation Types We observe that problems involving division and a higher number of operations tend to be more challenging for most LLMs (see Figure 8 in the Appendix for visualizations). For instance, problems requiring multiple division operations or a combination of different operations (e.g., addition, subtraction, multiplication, and division) are more likely to result in incorrect predictions. This aligns with the intuition that complex problems requiring more mathematical operations are generally more difficult for LLMs to solve (Imani et al., 2023b).
  • Tokenization Choices and Numerical Reasoning: Recent research has also highlighted the impact of tokenization choices on LLMs’ numerical reasoning capabilities (Singh and Strouse, 2024). Our experiments reveal similar patterns, with LLama-based models tokenizing each digit individually, while GPT-3.5 and GPT-4 encode every three digits into a single token. This difference in tokenization may contribute to the observed consistent performance of GPT-3.5 to the number of tokens in the answer compared to the LLama-based models.
Why LLMs suffer From Such Simple Attacks?
@@
While LLMs can be trained on vast amounts of math data, they may not have an inherent grasp of the underlying mathematical concepts and reasoning steps (Saxton et al., 2019; Lample and Charton, 2019). Instead, they likely rely on pattern and statistical associations learned from the training data to make predictions (Bender et al., 2021), which is also know as memorization (Carlini et al., 2022). Modifying the numbers in a problem may disrupt these learned patterns, causing the models to make errors. This is evidenced by the significant performance drops observed in models like WizardMath (89% to 20%) and Vicuna (76% to 4%) when presented with adversarial examples (Table 1). Furthermore, test data contamination could also play a role - if the training corpus contains the test data, the models may be able to answer the original questions correctly but struggle to generalize to even subtle variations (Balloccu et al., 2024).
Generated Python Code and ASTs Analysis
To ensure that the generated Python code is valid and can be correctly converted into ASTs, we manually examined 200 Python code and ASTs pairs from GSM8K and MultiArith, respectively. We identify the four types of errors:
  • Incorrect answers or code: The generated code produces incorrect answers or contains logical errors that deviate from the problem statement.
  • Use of unsupported constructs: The code includes loops, conditional and comparison statements, or user-defined functions, which are not supported by our AST conversion process.
  • Complex expressions: The code contains expressions that are difficult to convert into ASTs, such as x = y // z + (y % z > 0).
  • Number misalignment: The same number appears multiple times in the code or the problem, leading to inconsistencies between the generated code and the problem statement.
Encouragingly, we find that 96.5% and 92.5% of the generated ASTs for GSM8K and MultiArith, respectively, are valid and free from these errors. We discuss this process in more details in the Appendix H. The high percentage of valid code and ASTs pair indicates that our code generation approach is indeed reliable.
6 Conclusion
@@
We introduce a novel method to generate adversarial MWPs using ASTs. Our approach effectively challenges the mathematical problem-solving abilities of LLMs while maintaining the original difficulty and coherence of the problems. The generated adversarial examples significantly degrade the performance of both open- and closed-source LLMs, surpassing previous attacks by an average of 62% ASR. We validate our adversarial examples through human evaluation and investigate universal attacks and transferability, proposing a cost-effective method to attack high-cost API-based models with up to a 90% reduction in requests. Automatic regression analyses reveal distinct weaknesses of different models when solving MWPs. Our work contributes to the development of fair and robust educational tools, ensuring ethically sound evaluations and promoting the responsible use of LLMs in education.
7 Limitations
This work has not empirically validated the correlation between the complexity of generated problems and the actual difficulty perceived by human students. Further investigation is needed to establish this correlation.
As a proof of concept, our method demonstrates the feasibility of generating meaningful and useful adversarial examples to LLMs, at least at the difficulty levels represented by the GSM8K and MultiArith datasets. However, it may not generalize well to math problems with significantly different types or difficulty levels. For example, very simple arithmetic problems (e.g., 3 + 2 = ?) might never have a successful attack on LLMs specifically trained for math, while advanced problems requiring proofs or concepts without numeric manipulation (e.g., prove newton’s binomial theorem) may not be suitable for our code generation procedure. As LLMs continue to improve, it is possible that they may become more robust, which is a challenge faced by all adversarial attack methods in general. We plan to address these limitations in future work by exploring a wider range of math problems and investigating other attack vectors as LLMs evolve.
@@ 8 Ethics Statement
9 Acknowledgements
We are very grateful to Sandy Yeung and Jack Goffinet for the helpful initial project discussion. Thank you to the anonymous reviewers for their feedback. The research reported here was supported by the Learning Engineering Virtual Institute funded by leading education philanthropists and organizations through Grant G-23-2137070, to the University of Florida and its partner institutions. The opinions expressed are those of the authors and do not represent the views of the University of Florida, the partner institutions, or those of the philanthropists and organizations.
@@ A.1 Code Generation Prompt
Write a Python script that contains a step-by-step solution for a given math problem and its answer. Start the script by defining variables corresponding to the quantities mentioned in the problem. Print the final answer at the end of the script. Do not use if-else statements, floor divisions, and loops to solve the problem. Given problem: {problem} Answer: {answer} Python script:
@@ A.2 Zero-shot CoT
Solve a math problem. The solution ends with “the answer is (a number)" like “the answer is 1". Question: {problem} Answer: Let’s think step by step.
Appendix B A Rephrasing Attack Example
@@ Appendix C AST Nodes
We associate the number in the original question with the variable node so that a change in the variable node will also change the number in the question. If a number appears multiple times in the code and the problem, then we associate them based on the specific procedures described in Section H.4.
@@ Appendix E Rephrasing Attack Performance
Table 7: While rephrasing the problem results in lower accuracy for 4 out of 7 models, a few models show improved accuracy with the rephrased version. This suggests that rephrasing the MWPs may not generalize effectively across different LLMs.
@@ Appendix F Incorrect Variant Count Distribution
Figure 5: We present the distribution of incorrect adversarial examples in different buckets. For strong models (e.g., GPT-3.5-Turbo, MetaMath 70B), around half of the problems have zero incorrect adversarial examples, while for weaker models (e.g., Llama-2 13B, CodeLlama 34B), most problems have more than 90 incorrect adversarial examples.
@@ Appendix G Number of Generation
Figure 6: With 30,000 attempts for each given constraint, we calculate the average number of adversarial examples generated for each problem.
Appendix H Code Analysis
H.1 Incorrect Answer from the Dataset
Given the execution steps in code are almost always deterministic, we are able to identify this error effectively. There are also instances where the dataset-provided answers might be wrong. An example of such an instance is the following question from the MultiArith dataset: "Paige’s team won their dodgeball game and scored 41 points total. If Paige scored 11 of the points and everyone else scored 6 points each, how many players were on her team?", which the dataset provided 5 as the answer. However, the correct answer should be 6, as Paige is on the team as well. The code provided by GPT-4 counted Paige to the team. To automatically detect such instances, the value of the final answer node will be compared to the answer given by the dataset, and the question will be filtered out for further generations of problems if the two answers don’t match. In total, we found only 1% and 2.5% generated codes have incorrect logic or answers in GSM8K and MultiArith, respectively.
H.2 Incorrect Answer from GPT-4
Similarly, there are also instances where GPT-4 generated codes that provided the wrong answer to the given question. One such question is from the GSM8K dataset: "Jen decides to travel to 3 different countries. He has to pay \$400 for the supplies he needs, in total. The tickets for travel cost, in total, 50% more than the supplies. How much does travel cost?" The correct answer should be $400+400\times 1.5=1000$, but the code generated by GPT-4 outputted \$600. Similar to the previous error, instances where the code outputs the wrong answers will be detected by comparing them to answers given by the dataset, and the instances will be discarded from generating math problems.
H.3 Contains Control Flow Statements
Although specified in the prompt, GPT-4 might generate codes that include control flow statements (including "if", "for", and "while" statements). Codes that contain control flow statements are considered errors and discarded because the codes are usually only applicable to the original instance of the question. Furthermore, the code might not halt for some combinations of numbers, resulting in no answers associated with the generated instances.
H.4 Number Misalignment